OpenStudy (anonymous):

Solve for x: |3x + 2| = 2

OpenStudy (anonymous):

make two equations out of it, one positive and one negative.

OpenStudy (anonymous):

\[-3x+2=2\]

OpenStudy (anonymous):

\[3x+2=2\]

OpenStudy (hiroromon):

you are solving for x not getting rid of the x

OpenStudy (anonymous):

Thats how you solve...

OpenStudy (anonymous):

They are two separate equations smart one.

OpenStudy (hiroromon):

no you would get x on on side by subtracting the 2on the side with the x with the 2 on the other side

OpenStudy (anonymous):

Cause its absolute value so you must make two different equations and solve for x on both.

OpenStudy (hiroromon):

ok if you do it your way you would get 0 + 4 = 4

OpenStudy (anonymous):

No you wouldnt...I never did say to add them together? just said make two equations.

OpenStudy (anonymous):

why dont you go look up what absolute value is :)

OpenStudy (hiroromon):

>.>

OpenStudy (hiroromon):

anyway this is what you would do

OpenStudy (hiroromon):

the product of an abosulte value can be either negative or positive

OpenStudy (anonymous):

Right...

OpenStudy (hiroromon):

therefore in order to find the value of x you would normally solve it by doing this

OpenStudy (anonymous):

My way would be...\[-3x+2=2 -3x=0\]

OpenStudy (anonymous):

The problem is not answerable.

OpenStudy (hiroromon):

3x+2 = 2 -2 -2 = the answer below 3x=0 -- -- = undefined 3 3

OpenStudy (hiroromon):

and since it's absolute value you change the answer to a negative like so 3x+2 =-2

OpenStudy (hiroromon):

and you do the same thing

OpenStudy (anonymous):

But thats not how you do it in the future, for this problem that works but usually you must make two separate equations.

OpenStudy (anonymous):

Actually your wrong the answer is 0 and -3/4

OpenStudy (anonymous):

*-4/3